OpenStudy (anonymous):

The product of 8 and the sum of 4 and a number is 112. What is the number? Which equation could be used to solve for the number? 8 + 4n = 112 8 x 4 + n = 112 8(4 + n) = 112

OpenStudy (anonymous):

the top one???

OpenStudy (anonymous):

I believe it's the second one but wait for another answer just to be sure. :D

OpenStudy (anonymous):

No its the third one sorry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Because the top would have to use P.E.M.D.A.S. and it would be wrong so the last one

OpenStudy (anonymous):

the bottom would use pemdas to tho

OpenStudy (anonymous):

Yeh the paraentheses first thats what you want

OpenStudy (anonymous):

oh kk

OpenStudy (anonymous):

because you aren't doing 8 times 4 + x its 4 + x times 8 pretty much

OpenStudy (nirmalnema):

the correct answer is the 3rd eq... according to the ques....8 is multiplied to the sum of 4 and that unkonwn no..

OpenStudy (anonymous):

So the third one!!! :D LoL

OpenStudy (anonymous):

8(4+n) = 112 4+n=112/8 4+n=14 n=14-4 n=10 So its 8(4+n) = 112

OpenStudy (anonymous):

I hope I'm clear :)